OpenStudy (anonymous):

An awards dinner costs $225 plus $5 for each person making reservations. The total bill is $735. How many people made reservations?

OpenStudy (mertsj):

225+5x=735

Directrix (directrix):

Let P = number of persons making reservations. Let C represent the total cost of the bill C = 225 + 5P 735 = 225 + 5P 510 = 5P P = 102 --> number of persons making reservations